# U2A: Unified Unimodal Adaptation for Robust and Efficient Multimodal Learning

### v1

@@ Abstract
Multimodal learning often relies on designing new models and complex training strategies to achieve optimal performance. We present Unified Unimodal Adaptation (U2A), which jointly fine-tunes pretrained unimodal encoders using low-rank adaptation (LoRA) for various multimodal tasks. Our method significantly reduces the number of learnable parameters and eliminates the need for complex training strategies, such as alternating training, gradient modifications, or unimodal fine-tuning. To address missing modalities during both training and testing, we introduce Mask Tokens (MT), which generate missing modality features from available modalities using a single token per modality. This simplifies the process, removing the need for specialized feature estimation or prompt-tuning methods. Our evaluation demonstrates that U2A matches or outperforms state-of-the-art methods in both complete and missing modality settings, showcasing strong performance and robustness across various modalities, tasks, and datasets. We also analyze and report the effectiveness of Mask Tokens in different missing modality scenarios. Overall, our method provides a robust, flexible, and efficient solution for multimodal learning, with minimal computational overhead.
@@ 1 Introduction
Multimodal learning [1, 2] combines information from different input sources to enhance performance on downstream tasks. A majority of existing methods assume well-curated datasets in which all modalities are available for every sample. In real-world applications, input modalities can be missing during training and/or testing, and both cases can severely degrade model performance [3]. The modalities can be missing due to sensor failure, privacy concerns, or lack of synchronous data capture. In this paper, we propose a method that maintains strong performance in both complete and missing modality scenarios with limited computational overhead.
@@
Figure 1: Illustration of our proposed method that can accommodate complete and missing modalities during training and testing. We adapt pretrained unimodal encoders for each modality and introduce mask tokens to estimate missing modality features from available modalities. Our approach enhances robustness across various missing modality scenarios.
To integrate information from different input sources, existing methods focus on early-stage fusion of modalities [4], token-level fusion [5], channel exchange [6], and bottleneck mid-fusion [7]. Additionally, specialized architectures have been proposed for specific multimodal tasks, such as vision-language tasks [8, 9, 10], action recognition [11, 12], and segmentation [13, 14]. Designing specialized, task-specific architectures for every multimodal scenario is impractical due to the vast range of potential tasks. Instead, a preferred approach is to develop models and fusion strategies that can maintain performance and efficiency while being adaptable to diverse settings.
Models trained with multiple input modalities exhibit a drastic performance drop when one or more modalities are missing [3]. Several approaches have been proposed to enhance robustness against missing modalities. These include robust training techniques [15, 16], knowledge distillation methods [17, 18, 19], modality masking strategies [20, 21], generative networks [22, 23] to impute missing modalities, and multimodal prompting [24, 25, 26]. While effective, these methods involve learning a large number of parameters or prompts across multiple layers of the network, which increases computational and memory complexity.
In this paper, we propose Unified Unimodal Adaptation (U2A), a method that leverages pretrained unimodal encoders for multimodal learning. Our method encodes each modality separately through pretrained transformer encoders. To handle missing modalities, we introduce mask tokens to estimate the missing modality features from available input modalities. We introduce a modality alignment loss to learn the mask tokens that can effectively substitute for the class tokens of the missing modality. Our approach jointly optimizes this alignment loss with task-specific objectives through random modality dropout and low-rank adaptation (LoRA [27]) to learn cross-modal relationships while keeping computational costs low. Our method requires significantly fewer learnable parameters compared to state-of-the-art (SOTA) methods, while consistently delivering superior performance. We performed extensive experiments on five datasets, comparing our approach to existing baseline methods. The results show that our method outperforms SOTA methods in both complete (Section 4.3) and missing modality (Section 4.4) scenarios while requiring a small number of learnable parameters. Ablation studies confirm that our method can accurately estimate missing modality features, leading to improved performance in missing modality scenarios (Section 4.5). The main contributions of our paper can be summarized as follows.
  • We introduce Unified Unimodal Adaptation (U2A), which outperforms most of the existing methods in various multimodal tasks, while requiring fewer learnable parameters (Section 4.3).
  • We propose to learn Mask Tokens, which efficiently represent cross-modal relationships and infer missing modality features from available modalities, improving robustness in missing modality scenarios (Section 4.4).
  • We present extensive experiments that demonstrate our approach outperforms existing methods across multiple datasets and achieves strong performance in both complete and missing modality scenarios with fewer learnable parameters.
@@ 2 Related Work
2.1 Multimodal Learning
Designing specialized models to fuse multiple input sources effectively is a common practice in multimodal learning. For vision-language tasks, models like ViLT [8], ViLBERT [10], and Visual-BERT [9] process image patches and text tokens directly within a transformer model for tasks like image captioning and visual question answering. MMBT [28] integrates text and images within a BERT-based structure [29] to model cross-modal dependencies. MBT [30] uses fusion bottlenecks for efficient modality-specific interaction for audio-video tasks. A number of multimodal models are also proposed for other tasks like action recognition [11, 12], segmentation [13, 5] and sentiment analysis [11, 31]. While effective for specific applications, extending these models to new tasks or modalities remains non-trivial and resource-intensive.
Adapting existing models is another approach for multimodal learning. One such approach is gradient modulation. G-Blend [32] modulates gradients to combine unimodal and multimodal knowledge. OGM-GE [33] refines this method by emphasizing on unique information from each modality. PMR [34] modulates parameter groups based on modality relevance, while QMF [35] applies quadratic modulation for adaptive fusion. More recently, MLA [36] proposed alternating adaptation, which switches between unimodal adaptation and gradient modulation, enabling effective cross-modal learning. MMLoRA [37], on the other hand, uses unimodal fine-tuning followed by multimodal adaptation to achieve optimal performance. However, these approaches often require learning a large number of parameters to achieve optimal performance on downstream tasks.
2.2 Missing Modality Robustness
@@
Recent works [3, 38] show that performance can degrade significantly when modalities are missing, as multimodal models are not inherently robust to such scenarios. To address this issue, several approaches have been proposed.
@@
Robust model design is one of the approaches for handling missing modalities. Recently, Shin et al. [21] designed a robust framework utilizing modality masking and knowledge distillation. Wang et al. [39] proposed ShaSpec to impute missing features through modality-specific and shared feature learning. Wang et al. [40] proposed TokenFusion, which dynamically replaces uninformative tokens to enhance robustness. Additionally, general-purpose fusion strategies [41, 42, 43] aim to improve robustness across various missing-modality scenarios. However, many of these approaches are task-specific, limiting their generalization to other multimodal tasks.
@@
Robust training and model adaptation can enhance missing modality robustness. Recently Reza et al. [44] proposed a parameter-efficient adaptation framework for different multimodal tasks. Nezakati et al. [45] trained a single model to infer missing information by projecting available modalities. Ma et al. [46] introduced SMIL, a Bayesian meta-learning method for flexible handling of missing modalities during both training and testing. Ma et al. [3] proposed a multi-task optimization approach with a differentiable algorithm for optimal fusion strategy selection. Maheshwari et al. [18] developed a semi-supervised framework that leverages unlabeled data to improve robustness, while Wei et al. [47] used a teacher network to transfer multimodal information for improved missing modality robustness. These methods, however, either require extensive adaptation or specialized training procedures.
@@
Prompt tuning methods are also used to compensate for missing modalities. Lee et al. [24] introduced missing-aware prompts that learns a separate set of prompts for each modality combination. Later, Jang et al. [25] simplified the approach, showing that a single set of prompts per modality can achieve similar performance. More recently, Dai et al. [26] proposed a multi-step adaptive prompt learning method through modality alignment, while Kim and Kim [48] introduced a read-only prompt tuning approach. Although effective, these methods require learning a large number of prompts for each missing scenario.
In contrast, our method utilizes pretrained unimodal encoders, fine-tuning them jointly on multimodal data. To handle missing modalities, we introduce mask tokens that estimate missing modality features from the available ones. A modality alignment loss is jointly optimized with task-specific objectives using random modality dropout and LoRA to effectively learn cross-modal relationships. This design ensures strong performance in both complete and missing modality scenarios while offering the flexibility to add new modalities with minimal computational cost. Our approach is efficient, uses few learnable parameters, and is robust to missing modalities in both training and testing.
3 U2A: Unified Unimodal Adaptation
Our proposed unified unimodal adaptation (U2A) approach can accommodate complete and missing modalities during training and testing. Our goal is to develop a simple and efficient learning framework that can be used for different multimodal tasks without requiring specialized (task-dependent) changes. In this vein, we propose to adapt pretrained unimodal encoders for each modality and introduce mask tokens to estimate missing modality features from available modalities. Figure 1 illustrates main components of our approach that we describe in detail below.
@@ 3 U2A: Unified Unimodal Adaptation
Multimodal data embedding. We consider a multimodal dataset containing two modalities =$\mathcal{M}=\{m_{1},m_{2}\}$. Let us denote the available dataset as $\mathcal{D}=\{\mathcal{D}_{c},\mathcal{D}_{m_{1}},\mathcal{D}_{m_{2}}\}$, which can be further divided into three subsets: $\mathcal{D}_{c}=\{x_{m_{1}},x_{m_{2}},y\}$ with inputs from both modalities $(x_{m_{1}},x_{m_{2}})$ and output labels $y$; $\mathcal{D}_{m_{1}}=\{x_{m_{1}},\varnothing,y\}$ with inputs only from modality $m_{1}$ $(x_{m_{1}})$ and labels $y$; $\mathcal{D}_{m_{2}}=\{\varnothing,x_{m_{2}},y\}$ with inputs only from modality $m_{2}$ $(x_{m_{2}})$ and output labels $y$. $\varnothing$ denotes a placeholder for missing modality (e.g., empty string for texts, zeros for images, audios, and videos) [24, 44].
@@
Frozen embedding layers. For each input modality, we assume that a transformer-based encoder pretrained on some large dataset is available. Specifically, we use a pretrained modality-specific $\text{EmbeddingLayer}_{m}$ with parameters $\Theta_{e,m}$. Each input modality $x_{m}$ passes through the corresponding embedding layer to generate modality specific tokens as
@@
Adaptation of unimodal encoders. For each modality, we assume that a pretrained transformer encoder $f_{m}$ with parameters $\Theta_{f,m}$ is available. To adapt each encoder, we insert LoRA [27] layers with learnable parameters $\Delta_{f,m}$. We concatenate one class token $\mathcal{T}_{\text{CLS},m}\in\mathbb{R}^{1\times d}$ and one Mask Token (MT) $\mathcal{T}_{\text{MT},m}\in\mathbb{R}^{1\times d}$ with $\mathcal{T}_{m}$ to generate the updated tokens as
$\tilde{\mathcal{T}}_{m}=\text{Concat}(\{\mathcal{T}_{\text{CLS},m},\mathcal{T}_{m},\mathcal{T}_{\text{MT},m}\}),$ (2)
where $\tilde{\mathcal{T}}_{m}\in\mathbb{R}^{(N+2)\times d}$. The updated tokens are passed through the corresponding modality specific encoder $f_{m}$ to extract the feature tokens as
@@
where $\hat{\mathcal{T}}_{m}\in\mathbb{R}^{(N+2)\times d}$ represents the tokens/hidden-states from the last encoder layer. During training, we keep the pretrained parameters $\Theta_{f,m}$ frozen and only update the small number of learnable parameters $\Delta_{f,m}$.
Token selection. After extracting the feature tokens $\hat{\mathcal{T}}_{m}$, we take the class token $\hat{\mathcal{T}}_{\text{CLS},m}$ and the mask token $\hat{\mathcal{T}}_{\text{MT},m}$ for both modalities and pass them to a token selection block. If both modalities are available, the token selection block returns the class tokens from each modality. If any modality is missing, the block replaces the class token of the missing modality with the mask token from the available modality. We can define the selected tokens $\mathcal{T}_{1},\mathcal{T}_{2}\in\mathbb{R}^{1\times d}$ as
$(\mathcal{T}_{1},\mathcal{T}_{2})=\begin{cases}(\hat{\mathcal{T}}_{\text{CLS},{m_{1}}},\hat{\mathcal{T}}_{\text{CLS},{m_{2}}}),&m_{1},m_{2}\text{ are available,}\\ (\hat{\mathcal{T}}_{\text{MT},{m_{2}}},\hat{\mathcal{T}}_{\text{CLS},{m_{2}}}),&m_{1}\text{ is missing,}\\ (\hat{\mathcal{T}}_{\text{CLS},{m_{1}}},\hat{\mathcal{T}}_{\text{MT},{m_{1}}}),&m_{2}\text{ is missing.}\end{cases}$ (4)
Fusion and predictions. Finally, we sum up the selected tokens to generate the fused feature and pass it through the linear classifier head to make prediction $\hat{y}$ as
$\hat{y}=h(\mathcal{T}_{1}+\mathcal{T}_{2};\Theta_{h}).$ (5)
$\Theta_{h}$ denotes all the parameters of the classifier head that we learn during training.
Mask tokens alignment. To accurately estimate the feature of a missing modality using the mask token of the available modality, we introduce a modality alignment loss function. Specifically, for any sample $x\in\mathcal{D}_{c}$, our goal is to ensure that $\hat{\mathcal{T}}_{\text{MT},{m_{1}}}\approx\hat{\mathcal{T}}_{\text{CLS},{m_{2}}}$ and $\hat{\mathcal{T}}_{\text{MT},{m_{2}}}\approx\hat{\mathcal{T}}_{\text{CLS},{m_{1}}}$. We define the alignment loss as
$\mathcal{L}_{\text{align}}=\frac{1}{N_{\mathcal{D}_{c}}}\sum_{x\in\mathcal{D}_{c}}\begin{array}[]{l}\mathcal{L}_{\text{MSE}}(\hat{\mathcal{T}}_{\text{MT},{m_{1}}},\hat{\mathcal{T}}_{\text{CLS},{m_{2}}})\,+\\ \mathcal{L}_{\text{MSE}}(\hat{\mathcal{T}}_{\text{MT},{m_{2}}},\hat{\mathcal{T}}_{\text{CLS},{m_{1}}})\end{array}$ (6)
where ${N_{\mathcal{D}_{c}}}$ is the number of samples in modality-complete subset $\mathcal{D}_{c}$ and $\mathcal{L}_{\text{MSE}}$ denotes mean squared error loss.
Training loss. We use cross-entropy loss for all the classification tasks, which we denote as $\mathcal{L}_{\text{task}}$. The overall objective to minimize is

### v2

# Robust Multimodal Learning via Cross-Modal Proxy Tokens

@@ Abstract
Multimodal models often experience a significant performance drop when one or more modalities are missing during inference. To address this challenge, we propose a simple yet effective approach that enhances robustness to missing modalities while maintaining strong performance when all modalities are available. Our method introduces cross-modal proxy tokens (CMPTs), which approximate the class token of a missing modality by attending only to the tokens of the available modality. To efficiently learn the approximation for the missing modality via CMPTs with minimal computational overhead, we employ low-rank adapters in frozen unimodal encoders and jointly optimize an alignment loss with a task-specific loss. Extensive experiments on five multimodal datasets show that our method outperforms state-of-the-art baselines across various missing rates while achieving competitive results in complete-modality settings. Overall, our method offers a flexible and efficient solution for robust multimodal learning. The code and pretrained models will be released on GitHub.
@@ 1 Introduction
Figure 1: Our proposed approach, cross-modal proxy tokens (CMPTs), achieves state-of-the-art performance in both complete and missing modality scenarios. The radar plot shows F1-macro scores on the MM-IMDb dataset, with each axis representing the % availability of image and text modalities. We compare CMPTs against four recent methods [21, 29, 22, 20] using the same experimental setup as  [22, 20].
@@
Figure 2: Illustration of our proposed method. We use pretrained unimodal encoders and adapt them to learn cross-modal proxy tokens (CMPTs) that can approximate missing modality class tokens. We learn the CMPTs and adaptation parameters by jointly optimizing alignment and task-specific objectives. Our approach can accommodate complete and missing modalities during both training and inference and enhances robustness across various missing modality scenarios.
Multimodal learning [35, 52] integrates information from multiple (diverse) input sources to improve the performance on downstream tasks. For example, examining a movie poster image along with a short text synopsis can provide a greater insight into the movie’s type and genre. In recent years, a number of methods have been proposed to effectively combine information from multiple modalities, including early-stage fusion [31], token-level fusion [49], channel exchange [47], and bottleneck mid-fusion [32]. Furthermore, specialized architectures have been proposed for specific multimodal applications, such as vision-language tasks [21, 23, 27], action recognition [51, 56], and segmentation [53, 40]. Most of these approaches assume that all modalities are available for every sample during both training and inference, yielding better performance than unimodal baselines. However, their performance drops sharply when one or more modalities get missing [39, 29].
In practical applications, input modalities can be missing during training and/or inference for various reasons; such as lack of synchronous data capture (limited paired data during training) and sensor failure or privacy concerns (modalities missing during inference). Ma et al. [29] observed that multimodal transformers such as ViLT [21] exhibit a drastic performance drop when one or more modalities are missing during inference as shown in Figure 1 (yellow curve). They developed an optimal fusion strategy to recover the performance loss of multimodal transformers in the missing modality scenarios (purple curve). Subsequently, Lee et al. [22] developed a prompt tuning technique (pink and blue curves), and Kim et al. [20] combined the prompt-based approach with Variance Covariance Invariance regularization (VICReg) (brown curve) to further improve the robustness of multimodal models to missing modalities. Achieving high performance when modalities are missing at varying rates remains a challenge, which we seek to address in this paper.
In this paper, we introduce a simple yet effective framework for multimodal learning that leverages pre-trained unimodal encoders to approximate missing modalities from the available ones. Specifically, we introduce cross-modal proxy tokens (CMPTs) that learn an approximation of the class tokens of the missing modality by attending on all the tokens of the available modality to handle different missing scenarios. As illustrated in Figure 2, our framework integrates pre-trained unimodal encoders into a multimodal learning setting while introducing CMPTs to bridge the performance gap in missing modality scenarios. To ensure efficient learning of CMPTs with minimal additional learnable parameters, we employ low-rank adapters within the frozen unimodal encoders and jointly optimize both an alignment loss and a task-specific loss. The radar plot in Figure 1 demonstrates the superior performance of our approach compared to five existing methods [21, 22, 20, 29] across six different complete and missing modality scenarios on the MM-IMDb dataset, where our CMPT-based model consistently outperforms all baselines. We conduct extensive experiments on five multimodal datasets, comparing our approach against existing state-of-the-art baselines. Results show that our framework achieves state-of-the-art performance in both missing modality (Section 4.3) and complete modality (Section 4.4) scenarios, while maintaining a minimal number of learnable parameters. Ablation studies further confirm that our method significantly improves performance even in severe missing-modality scenarios and provides performance enhancement across most of the classes, leading to an overall performance boost (Section 4.6).
Main contributions of our work are as follows.
  • We propose cross-modal proxy tokens (CMPTs) that learn an approximation of the missing modality from the available modality with minimal adaption of pre-train unimodal encoders via an alignment loss.
  • Our approach outperforms 12 recent baseline methods across four multimodal datasets in various missing-modality scenarios (Section 4.3).
  • Across five datasets, our method achieves on-par or superior performance compared to state-of-the-art techniques when all modalities are present during training and inference (Section 4.4).
  • We demonstrate that CMPTs generalize effectively to unseen missing rates (Section 4.5) and boost performance in most of the classes (Section 4.6).
@@ 2 Related Work
Recent works [29, 14, 39] show that performance can degrade significantly when modalities are missing, as multimodal models are not inherently robust to such scenarios. Several approaches have been proposed to address this issue. In this section, we primarily discuss works where the main focus is on making multimodal models robust to missing modalities. More broader related work on multimodal learning is discussed in Supplementary Section S1.
@@
Robust model design is one of the approaches for handling missing modalities. Recently, Shin et al. [41] developed a robust framework utilizing modality masking and knowledge distillation. Wang et al. [44] proposed ShaSpec to impute missing features through modality-specific and modality-shared feature learning. Wang et al. [48] proposed TokenFusion, which dynamically replaces uninformative tokens to enhance robustness. Additionally, different fusion strategies [5, 10, 25] have been developed to improve robustness across various missing-modality scenarios. However, many of these approaches are task-specific, limiting their generalization to other multimodal tasks. In our work, we focus on reusing the pre-trained unimodal encoders, making it flexible for extension to different input modalities and multimodal tasks.
@@
Robust training and model adaptation can enhance missing modality robustness. Recently Reza et al. [39] proposed a parameter-efficient adaptation framework for different multimodal tasks. Nezakati et al. [34] trained a single model to infer missing information by projecting available modalities. Ma et al. [28] introduced SMIL, a Bayesian meta-learning method for flexible handling of missing modalities during both training and testing. Ma et al. [29] proposed a multi-task optimization approach with a differentiable algorithm for optimal fusion strategy selection. Maheshwari et al. [30] developed a semi-supervised framework that leverages unlabeled data to improve robustness, while Wei et al. [50] used a teacher network to transfer multimodal information for improved missing modality robustness. These methods, however, either require extensive adaptation or specialized training procedures. In our approach, we employ lightweight adaptation of the unimodal encoders (using Rank-1 LoRA adapters) and a single extra loss term to explicitly learn cross-modal relationships. Hence, the overall training flow is mostly left unchanged.
@@
Prompt tuning methods improve missing modality performance by incorporating learnable prompts across multiple layers of the model. Lee et al. [22] introduced missing-aware prompts that learns a separate set of prompts for each modality combination. Later, Jang et al. [16] simplified the approach, showing that a single set of prompts per modality can achieve similar performance. More recently, Dai et al. [6] proposed a multi-step adaptive prompt learning method through modality alignment, while Kim and Kim [20] introduced a read-only prompt tuning approach. Although effective, these methods require learning and storing a large number of prompts, typically 16 prompts [22, 16] per modality per layer, for each missing modality scenario. In contrast, our method uses a single cross-modal proxy token per modality to estimate the missing modality feature from the available modality, thereby reducing the complexity. We carry out extensive experimental comparisons with all the above prompt-based techniques to demonstrate superior performance across various missing modality scenarios. Furthermore, our approach is efficient, uses few learnable parameters, and is robust to missing modalities in both training and testing which we discuss in Section 4.
3 Proposed Approach
In this paper, we propose a new framework to train multimodal systems such that they retain their performance under different scenarios of missing modalities during inference. Furthermore, we seek to develop a simple and efficient framework that can be used for different multimodal tasks and datasets without any specialized (task-specific) changes. To achieve these objectives, we propose to leverage modality-specific (transformer-based) encoders that are pre-trained on large datasets and align the features across modalities using proxy tokens, as illustrated in Figure 2.
3.1 Multimodal embedding tokens
@@
Let us consider a multimodal dataset containing two modalities =$\mathcal{M}=\{m_{1},m_{2}\}$. We denote the available dataset as $\mathcal{D}=\{\mathcal{D}_{c},\mathcal{D}_{m_{1}},\mathcal{D}_{m_{2}}\}$, which can be further divided into three subsets: $\mathcal{D}_{c}=\{x_{m_{1}},x_{m_{2}},y\}$ where both modalities are present with corresponding output labels $y$; $\mathcal{D}_{m_{1}}=\{x_{m_{1}},\varnothing,y\}$ where only modality $m_{1}$ $(x_{m_{1}})$ is available with corresponding labels $y$; $\mathcal{D}_{m_{2}}=\{\varnothing,x_{m_{2}},y\}$ where only modality $m_{2}$ $(x_{m_{2}})$ is available with corresponding labels $y$. Here, $\varnothing$ denotes a placeholder for the missing modality (e.g., empty string for text, zeros for image, audio, or video) following [22, 39].
@@
For each input modality, we assume that a transformer-based encoder, pre-trained on some large dataset, is available. Specifically, each modality $m\in\mathcal{M}$ is processed through a pre-trained embedding layer, $\text{EmbeddingLayer}_{m}$, with parameters $\Theta_{e,m}$. The input $x_{m}$ for each modality is passed through its corresponding embedding layer to generate modality-specific tokens as
@@
3.2 Cross-modal proxy tokens
To compensate for the missing modalities, we introduce cross-modal proxy tokens (CMPTs) that approximate the missing modality by attending only to the tokens of the available modality. For each modality $m\in\mathcal{M}$, we concatenate one CMPT ($\mathcal{T}_{\text{CMPT},m}\in\mathbb{R}^{1\times d}$) with the class token (CLS) ($\mathcal{T}_{\text{CLS},m}\in\mathbb{R}^{1\times d}$) and all the modality specific tokens $\mathcal{T}_{m}$ to generate the modified embedding tokens $\tilde{\mathcal{T}}_{m}\in\mathbb{R}^{(N+2)\times d}$ as
$\tilde{\mathcal{T}}_{m}=\text{Concat}(\{\mathcal{T}_{\text{CMPT},m},\mathcal{T}_{\text{CLS},m},\mathcal{T}_{m}\}).$ (2)
Note that the proposed CMPT serves as an extra token, similar to the class token [8, 7], but with a slightly orthogonal purpose. Class token attends to all the tokens of a given modality to aggregate information and learn a rich feature representation of that modality. In contrast, CMPT attends to all the tokens of a given modality to learn an approximation of the class token of the other modality (e.g., $\mathcal{T}_{\text{CMPT},m_{1}}$ approximates $\mathcal{T}_{\text{CLS},m_{2}}$ while attending to $\mathcal{T}_{m_{1}}$).
3.3 Learn CMPTs by alignment and adaptation
For each modality, we assume that a pretrained transformer encoder $f_{m}$ with parameters $\Theta_{f,m}$ is available. The encoder for each modality needs to be updated so that they can learn the corresponding CMPTs. To keep the number of trainable parameters minimal, we use low-rank adaptation (LoRA) [15] to learn a small number of parameters $\Delta_{f,m}$. The modified embedding tokens for modality $m$, $\tilde{\mathcal{T}}_{m}$, are then passed through the adapted encoder for modality $m$ to get the final output tokens as
@@ 3.3 Learn CMPTs by alignment and adaptation
where $\hat{\mathcal{T}}_{m}\in\mathbb{R}^{(N+2)\times d}$ represents the CMPT, CLS, and remaining tokens from the last encoder layer. Note that the pretrained encoder parameters $\Theta_{f,m}$ remain frozen while we learn $\Delta_{f,m}$.
We aim to train the CMPTs such that they act as a proxy of the class token of the missing modality at the inference time. In particular, we want the CMPT $\hat{\mathcal{T}}_{\text{CMPT},m_{1}}$ of modality $m_{1}$ to approximate the class token $\hat{\mathcal{T}}_{\text{CLS},m_{2}}$ of modality $m_{2}$, and vice versa. To achieve this, we use an alignment loss $\mathcal{L}_{\text{align}}$, defined as
$\mathcal{L}_{\text{align}}=\frac{1}{N_{\mathcal{D}_{c}}}\sum_{x\in\mathcal{D}_{c}}\begin{array}[]{l}\mathcal{L}_{\text{MSE}}(\hat{\mathcal{T}}_{\text{CMPT},{m_{1}}},\hat{\mathcal{T}}_{\text{CLS},{m_{2}}})\,+\\ \mathcal{L}_{\text{MSE}}(\hat{\mathcal{T}}_{\text{CMPT},{m_{2}}},\hat{\mathcal{T}}_{\text{CLS},{m_{1}}}),\end{array}$ (4)
where $\mathcal{L}_{\text{MSE}}$ denotes the mean squared error loss and ${N_{\mathcal{D}_{c}}}$ denotes the number of samples in the training data where both modalities are present.
3.4 Gating module and fusion
After computing the final output tokens $\hat{\mathcal{T}}_{m}$, we extract the class token ($\hat{\mathcal{T}}_{\text{CLS},m}$) and the CMPT ($\hat{\mathcal{T}}_{\text{CMPT},m}$) for both modalities and pass them through the gating module. If both modalities are available, the gating module returns the class tokens from each modality. If one modality is missing, the gating module substitutes the missing class token with the CMPT from the available modality. The returned tokens are then fused using additive fusion to produce the final fused feature token $\mathcal{T}\in\mathbb{R}^{1\times d}$ as
$\mathcal{T}=\begin{cases}\hat{\mathcal{T}}_{\text{CLS},{m_{1}}}+\hat{\mathcal{T}}_{\text{CLS},{m_{2}}},&m_{1},m_{2}\text{ are available,}\\ \hat{\mathcal{T}}_{\text{CLS},{m_{2}}}+\hat{\mathcal{T}}_{\text{CMPT},{m_{2}}},&m_{1}\text{ is missing,}\\ \hat{\mathcal{T}}_{\text{CLS},{m_{1}}}+\hat{\mathcal{T}}_{\text{CMPT},{m_{1}}},&m_{2}\text{ is missing.}\end{cases}$ (5)
The fused feature token $\mathcal{T}$ is then passed through a linear classifier head to make a prediction $\hat{y}$ as
$\hat{y}=h(\mathcal{T};\Theta_{h}).$ (6)
$\Theta_{h}$ represents the parameters of the classifier head that are learned during training.
Finally, we learn the low-rank factors for each encoder ($(\{\Delta_{f,m}\})$ and classifier head $(\Theta_{h})$ by minimizing the task-specific loss $\mathcal{L}_{\text{task}}$ combined with the alignment loss: